OpenStudy (anonymous):

the horsepower produced by a steam engine varies directly as the number of revolution per minute of its operation . if 1200 horse power was produced when a machine was operated at 150 revolutions per minute , find the number of horse power it will produce when it will operate at 170 revolution per minute . ? The current I in amperes is inversely proportional to resistance R in ohms . if the current is 30 amperes when the resistance is 3 ohms what is the current when the resistance is 5 ohms?

OpenStudy (anonymous):

1200 = 150 m y = 170 m

OpenStudy (anonymous):

second one, similar 300 = m/3 y = m/5

OpenStudy (anonymous):

you should just solve for m in both cases

OpenStudy (anonymous):

and then use that to find y

OpenStudy (anonymous):

what the is the formula that used ?

OpenStudy (anonymous):

Well, direct variation is \[y = mx\] where m is called the constant of variation. Inverse variation is \[y = \frac{m}{x}\] and m is also the constant of variation

OpenStudy (anonymous):

I used this approach because that is the way the question was phrased, you could also use a proportion, such as\[\frac{1200}{150}=\frac{y}{170}\]

OpenStudy (anonymous):

what is different of m and k ?

OpenStudy (anonymous):

you could use k if you wanted

OpenStudy (anonymous):

they're just variables

OpenStudy (anonymous):

ahhhh. ok .

OpenStudy (anonymous):

you could also do \[y = \beta x\] or  Variables are whatever you want! :)